Under its executive stock option plan, Muehlmann Corporation granted options on January 1, 2024, that permit executives to purchase 25.5 million of the company's $1 par common shares within the next eight years, but not before December 31, 2026 (the vesting date). The exercise price is the market price of the shares on the date of grant, $19 per share. The fair value of the options, estimated by an appropriate option pricing model, is $4 per option. No forfeitures were anticipated; however, unexpected turnover during 2025 caused the forfeiture of 5% of the stock options. Ignoring taxes, what is the effect on earnings in 2026?

Note: Round your answer to 1 decimal place
Answer #1

step 1: 25.5*4= 102        step 2) 102/3yrs= 34        2026-2024= 3 yrs

{102*95%(100%-5%)*2/3}-34= 30.6

 {102*95%(100%-5%)*3/3}-34-30.6= 32.3 answer

answered by: anonymous